OpenStudy (anonymous):

Sarah can paddle her kayak in still water with a speed of 20 meters per minute. if Sarah paddles 100 meters with the current in the same time she paddles 60 meters against the current. find the speed of the current i am having trouble finding the equation..the rest i can handle..plz help

OpenStudy (anonymous):

this is what i thought... you know in the formula for speed is speed= distance/time if u rearrange to make time the subject you get time=distance divided by speed....since she does the 100m and 60m in the same time then 100m/some speed = 60m/some speed...this is where i am lost...what to do from here...

OpenStudy (anonymous):

when she travels with the current ..her sped and the speed of current get sumed up ...but when going against ..her speed-the speed of current

OpenStudy (anonymous):

so.. let speed of current be x..then... 100/20+x=60/20-x?

OpenStudy (anonymous):

let the speed of the current be x m/minute so time taken while traveling 100m with the current =100/20+x and time while against 60/20-x equate these two and find x!

OpenStudy (anonymous):

u are right .!

OpenStudy (anonymous):

i did this...but i gotwrong answer..i did this way earlier today...wait..let me check..i think i did a lil mistake

OpenStudy (anonymous):

aha..its right..i did a stupid mistake..thx anyway though :D

OpenStudy (anonymous):

anytime .:)

OpenStudy (anonymous):

what i did wrong was this,..when cross multyplying.... on the left side..it was supposed to be 100*(20-x) what i did was 20-x(100) omg soo stupid..wasted an houur when it shouldve been 5 min lol thx

OpenStudy (anonymous):

ohh ..happens /! ..practice ..:)